Efficient Implementation of Sequential Loops in Dataflow Computation

Computation Structures Group Memo 350 June 5, 1993

Boon Seong Ang

Appeared in the Proceedings of the 1993 FPCA, Copenhagen, Denmark. Pages 169 - 178.

This paper describes research done at the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-89-J-1988.